# Positive-Unlabeled Constraint Learning (PUCL) for Inferring Nonlinear Continuous Constraints Functions from Expert Demonstrations

### v1

@@ III-C Iterative Learning Framework with Policy Filter and Memory Buffer
Last subsection discussed about learning constraint from a given unsafe set $\mathcal{P}$ as well as demonstrations $\mathcal{D}$. This and next subsection will elaborate on the generation of unsafe set $\mathcal{P}$. Existing methods generate $\mathcal{P}$ either in an iterative manner [8, 4, 9] or in a one-batch manner [3, 13, 7]. The former iteratively updates policy with current constraint function and generate new trajectories $\mathcal{P}$ to update the constraint. Since the policy is always updated with respect to up-to-date constraint network, the new trajectories generated in each iteration are mostly distributed around the boundary of the current constraint network, which will improve the accuracy of the constraint boundary. Therefore, this manner is efficient to learn the constraint boundary and can automatically adjust the distribution of the generated trajectories.
In contrast, the one-batch manner generates a bulk of trajectories $\mathcal{P}$ uniformly[3] or greedily[7] at the start of training. The following training only takes place on only existing $\mathcal{P}$, with no new data generated. This manner is more stable but relatively inefficient and poses a higher demand on the diversity and the coverage of the generated trajectories.
This work follows the iterative framework, and introduce a policy filter and a memory buffer both from our previous work [15]. Fig. 2 gives a sketch of the whole iterative structure.
Policy filter: In each iteration of policy updating, the policy is updated for fixed steps with the current constraint network $\zeta(s;\theta)$. To prevent bad updates from bad policies, a policy filter (6) is introduced to select trajectories with relatively high reward than the demonstration, which are believed to truly violate the unknown constraint according to Assumption 1.
$\mathcal{P}\leftarrow\{\tau_{i}\in\mathcal{P}|r(\tau_{i})\geq(1-\delta)r(\tau_{i}^{\mathcal{D}})\}$ (6)
@@
Memory buffer: Most iterative-framework-based papers generate new trajectories in each iteration and discard those from previous iterations [8, 4, 9]. Our previous work [15] identified that such a training manner tends to forget the infeasible region learned in early iterations. Thus, we introduce a memory replay buffer and record all the reliable infeasible data $\mathcal{R}$ identified in each iteration into the memory buffer $\mathcal{M}$. In the following iterations of learning constraints, both the current $\mathcal{R}$ and memory buffer $\mathcal{M}$ serve as infeasible data for training (see (5)).
@@ III-D Represent and Learn Policy via Constrained RL or Dynamical System Modulation
In the iterative framework discussed in the last subsection, we need to maintain a policy to generates high-reward trajectories while satisfying the currently learned constraints. We consider and compare two approaches: 1) constrained RL, which is effective at learning constrained policies in a high-dimensional environment with complex reward function but suffers from a slow and unstable training process[20], and 2) dynamical system modulation (DSM), which modulates a nominal policy with a rotation matrix to satisfy safety constraints. It features minimal training time and provides a controller with stability guarantee. This method is applicable and very suitable for robotic tasks where the control command is the velocity of each state (i.e., joint angle or end-effector pose).
@@
Constrained reinforcement learning: a networked policy is trained with a modern constrained RL algorithm PID-Lagrangian[20], which offers the advantage of automatically and more stably adjusting the penalty weight. As shown in (7), it reshapes the original reward by incorporating the constraint as a penalty term into the original reward function to avoid the infeasible states. Here, $w_{p}$ is a penalty weight adjusted by PID-Lagrangian updating rules [20, 21, 22], and $c(s)$ is the constraint indicator function.
Using this reshaped $r^{\prime}(s,a)$, a constrained optimal policy can be straightforwardly learned with the standard RL algorithm PPO [23].
$r^{\prime}(s,a)=r(s,a)-w_{p}c_{\theta}(s),\quad\text{where}$ (7)
$c_{\theta}(s)=\begin{cases}0&\text{if $\zeta_{\theta}(s)>d$ (feasible)}\\ 1&\text{if $\zeta_{\theta}(s)\leq d$ (infeasible)}\\ \end{cases}$
@@
Dynamical System Modulation: For the task where the control input is the velocity or acceleration of each state dimension, DSM can be applied to obtain a safe control policy by modulating nominal policy with the constraint network. The nominal dynamical system policy is represented as (8), where $s_{g}$ is the goal state and matrix $A(s)$ is a Gaussian Mixture Regression model learned from the known reward function following [24].
$\pi(s)=\dot{s}=A(s)(s-s_{g})$ (8)
@@
Given a learned constraint network $\zeta(s;\theta)$ whose output ranges from 0 to 1, one can obtain a provably safe policy $\pi(s;\zeta)$ by modulating the nominal policy following (9) [25], where $n(s)$ is the normal vector to the constraint boundary, $e_{i}(s)$ are the tangent vectors, $r(s)$ is the vector towards a reference point, $\lambda_{i}(s)$ modify the velocity along different vectors, and $\Gamma(s)$ specifies the relative distance of current state $s$ to the constraint boundary.
$\pi(s)=M(s)A(s)(s-s_{g}),\quad\text{where}$ (9a)
$M(s)=E(s)D(s)E^{-1}(s),$ (9b)
$E(s)=[r(s),e_{1}(s),e_{2}(s),...e_{n-1}(s)],$ (9c)
$n(s)=\frac{d\zeta(s)}{ds},n(s)\perp e_{1}(s)\perp e_{2}(s)\perp...e_{n-1}(s)$ (9d)
$D(s)=diag[\lambda_{1}(s),\lambda_{2}(s),\lambda_{3}(s)...,\lambda_{n}(s)],$ (9e)
$\lambda_{1}(s)=1-\frac{1}{\Gamma(s)},\lambda_{2,3,...,n}(s)=1+\frac{1}{\Gamma(s)},$ (9f)
$\Gamma(s)=1+10*(\zeta(s)-0.5)^{0.2}$ (9g)
Finally, the pseudo-code for the algorithm is given in Algorithm 1.
@@
1:  input: Demonstration set $\mathcal{D}$, randomly initialized constraint network $\zeta_{\theta}$, randomly initialized policy network $\pi_{\phi}$, empty memory buffer $\mathcal{M}=\varnothing$
@@
3:     Update the policy $\pi_{\phi}$ with the reshaped reward (7) and the current constraint network $\zeta_{\theta}$ using constrained RL or DS modulation
4:     Sample trajectory set $\mathcal{P}$ using the current policy $\pi_{\phi}$ from the same starting points as each demonstration
5:     Filter sampled set $\mathcal{P}$ with (6): $\mathcal{P}\leftarrow\{\tau_{i}\in\mathcal{P}|r(\tau_{i})\geq(1-\delta)r(\tau_{i}^{\mathcal{D}})\}$
6:     Identify and expand reliable infeasible set $\mathcal{R}$ using metric (3) and (4) =$\mathcal{R}=\{s\in\mathcal{P}|S_{\mathcal{D}}(s)\geq d_{r}\}$$\mathcal{R}\leftarrow\mathcal{R}\cup\{s|\exists\tau_{i}\in\mathcal{P},s=\arg\min_{s\in\tau_{i}}S_{\mathcal{R}}(s)\}$
7:     Train the constraint network $\zeta_{\theta}$ with the demonstration $\mathcal{D}$, reliable infeasible set $\mathcal{R}$ and memory buffer $\mathcal{M}$ $\mathcal{L}(\theta)=-\frac{1}{N+M}[\sum^{N}_{s_{i}\sim\mathcal{D}}\log\zeta_{\theta}(s_{i})-\sum^{M}_{s_{j}\sim\mathcal{R}\cup\mathcal{M}}\log(1-\zeta_{\theta}(s_{j}))]$
8:     Save all data from $\mathcal{R}$ into the memory buffer $\mathcal{M}\leftarrow\mathcal{M}\cup\mathcal{R}$
9:  until meet some stop criterion
@@ IV Experiment
IV-A Experiment Setup
Environment and constraint: to examine the performance of the proposed method, we apply it to learn two position constraints of the end-effector’s trajectory of a Panda robot arm, one in 2D and another in 3D spaces. In both tasks, the agent is initialized randomly and rewarded to reach a goal state. The state space consists of 2D or 3D positions of the end-effector and the control variables are their corresponding velocities. In the 2D reaching task, the true constraint is avoiding an irregular obstacle composed of 2 ellipses as shown in Fig. 3. In the 3D reaching task, the true constraint is avoiding tall cylinder(s).
We emphasize that learning these two constraints is nontrivial and challenging in our problem setting for several reasons. Firstly, we assume no prior knowledge of the obstacles’ size, shape, and location, nor of the true constraints, while some existing methods learn constraints of obstacles with known shape (i.e., a box or a cylinder) [13]. Secondly, there is no requirement for a differentiable environmental model (if using constrained RL to learn policy). Thirdly, the constraints are nonlinear, not just (the union of) boxes or hyperplanes, which contrasts with previous IRL-based constraint learning works that only learn a linear plain constraint such as $x\geq-3$ [8, 9].
The expert demonstration sets for two tasks consists of 4 or 10 safe trajectories, respectively. All demonstrations are generated by an entropy-regularized suboptimal RL agent, trained assuming full knowledge of the true constraint. The reward function in both tasks are the negative of the Cartesian path length to reach the goal, which is a natural choice for a goal-reaching task.
(a) PUCL (ours)
(b) MECL
Figure 3: True constraint and constraint learned with the proposed method PUCL and baseline MECL [8] in the 2D reaching task. The two white ellipses represent the boundary of the true constraint. The heat map is a visualization of the learned constraint $\zeta_{\theta}$, where the red region of $\zeta=0$ represents infeasibility, while the blue regions of $\zeta=1$ indicate feasibility. The trajectories of the demonstrations and the policy are shown in green and yellow, respectively. The red points correspond to identified reliable infeasible data $\mathcal{R}\cup\mathcal{M}$ (only for PUCL).
Baselines and implementations: Our baselines are chosen to study the following two questions:
  1. 1.
    How effective is our PUCL method compared with other constraint learning algorithms with similar problem setting?
  2. 2.
    In our PUCL framework, what are the differences between using constrained RL and dynamical system modulation to obtain the policy?
@@ IV-A Experiment Setup
To answer question 1, we compare the constraint learning performance of the four methods, all using constrained RL as the policy: 1) PUCL (the proposed method); 2) maximum-likelihood constraint learning (MECL) [8], a popular IRL-based method that can learn a continuous arbitrary constraint function; 3) binary classifier (BC) [8], a method directly trains an NN feasibility classifier using loss (5) but with all data from $\mathcal{P}$ as infeasible data; 4) generative-PUCL (GPUCL), an alternative PUCL approach using a generative model to identify reliable infeasible data. It first learns a Gaussian Mixture Model (GMM) of the expert distribution, and then identifies reliable infeasible data as the data from $\mathcal{P}$ with the lowest likelihood of being generated by the GMM. GPUCL is also first formulated in this paper but do not discussed in detail due to weaker performance. The results of comparison of the above four methods are presented in section IV-B.
To answer question 2, we conduct another experiment implementing our PUCL approach using 1) a constrained RL policy or 2) a DSM policy (see section III-D). The two frameworks are named CRL-PUCL and DSM-PUCL, respectively, and the results of comparison are presented in section IV-C.
@@
A list of important hyperparameters of our algorithm is given in Table I. The hidden activation functions in all networks are Leaky ReLU. The output activation function of the constraint network is the sigmoid function, whereas that of the policy and value network is tanh. Other baselines all adopt the same network architecture. Please refer to our code repo (released later) for other parameters.
TABLE I: List of hyperparameters of PUCL. For the neural network architectures, the number of hidden units in each layer is mentioned.
@@
PID-Lagrangian $K_{p}$ 1 20
PID-Lagrangian $K_{i}$ 0.1 0.5
$\mathcal{R}$ identification threshold $d_{r}$ 0.12 0.03
@@
Metrics: we present the evaluation utilizing two metrics: 1) the IoU (the intersection over union) index, which measures the correctness of the learned constraints. We uniformly sample points in the state space, and compute IoU as (Number of points which are both predicted to be infeasible and truly infeasible) divided by (Number of points either predicted to be infeasible or truly infeasible); 2) the unsafe rate, the per-step true constraint violation rate of the policy learned/modulated from the learned constraint.
(a) IoU - 2D task
(b) Unsafe rate - 2D task
(c) IoU - 3D task
(d) Unsafe rate - 3D task
Figure 4: The IoU index (higher is better) and unsafe rate (lower is better) of learned constraint function. The two figures on the left are for the 2D reaching task, while the two on the right are for the 3D reaching task. The x-axis indicates the training process, the number of timesteps the agent takes in the environment. Our method outperforms other baseline methods in both environments.
IV-B Results of Different Constraint Learning Methods
The IoU index and the unsafe rate in the two tasks are presented in Fig. 4. All the results are the average of 10 independent runs, and the shaded area represents the variance. Each group of 10 runs uses the same demonstrations and hyper-parameters, but differ in the initialization parameters for the networks.
@@ IV-B Results of Different Constraint Learning Methods
The proposed method PUCL exhibits superior performance compared to the baseline across both environments and metrics. It achieves a nearly zero unsafe rate at the convergence of the training in both environments. MECL suffers from performance decrease in the late training stage of the 2D task, this is partly caused by the constraint forgetting problem [15]. GPUCL adopts a similar framework to PUCL but performs worse than PUCL, possibly due to the difficulty of fitting a good expert distribution with a very limited numbers of demonstrations.
Despite the superiority of PUCL, one may question why the IoU metric of PUCL is still below 0.8 (2D task) or 0.6 (3D task), relatively far from 1. This is due to the fundamental difficulty of learning the nonlinear constraint without knowing the true parameterization of the constraint and with a limited number of demonstrations. There would be regions in the state space whose feasibility cannot be inferred and decided from the given demonstrations, thus restricting the learning accuracy.
@@
To provide readers with an intuitive understanding, we include visualizations of the 2D constraint learned with PUCL and MECL in Fig. 3. Comparing the learned constraints (red region) with the true constraints (white ellipses) confirms the effectiveness of our method in acquiring a model of a nonlinear constraint. Although the learned constraint area is partially incorrect due to limited demonstrations, it effectively captures the essence of the true constraint and is sufficient for training a safe policy. The baseline method MECL learns a constraint that is more inaccurate, and even some states visited by the demonstrations are classified as infeasible.
IV-C Results of Different Policy Representation and Learning Methods

### v2

@@ III-C Iterative Learning Framework with Policy Filter and Memory Buffer
Last subsection discussed learning constraint from a given unlabeled set $\mathcal{P}$ and demonstrations $\mathcal{D}$. This subsection and the next will elaborate on the generation of unlabeled set $\mathcal{P}$. Existing methods generate $\mathcal{P}$ either in an iterative manner [8, 4, 9] or in a one-batch manner [3, 13, 7]. The former iteratively updates policy with current constraint function and generates new trajectories to update the constraint. Since the policy is always updated with respect to the up-to-date constraint network, the new trajectories generated in each iteration are mostly distributed around the boundary of the current constraint network, which will improve the accuracy of the constraint boundary. Therefore, this manner is efficient to learn the constraint boundary and can automatically adjust the distribution of the generated trajectories. In contrast, the one-batch manner generates a bulk of trajectories uniformly[3] or greedily[7] at the start of training. The following training takes place only on those trajectories, with no new data generated. This manner is more stable but relatively inefficient and poses a higher demand on the diversity and the coverage of the generated trajectories.
This work follows the iterative framework, and introduces a policy filter and a memory buffer. Fig. 2 gives a sketch of the whole iterative structure.
Policy filter: In each iteration of policy updating, the policy is updated for fixed steps with the current constraint network $\zeta_{\theta}(\bar{s})$. To prevent undesired updates from poor policies, a policy filter (6) is introduced to select trajectories with a relatively high reward than the demonstration, which are believed to truly violate the unknown constraint according to Assumption 1.
$\mathcal{P}_{\tau}\leftarrow\{\tau_{i}\in\mathcal{P}_{\tau}|(1-\delta)r(\tau_{i})\geq r(\tau_{i}^{\mathcal{D}})\}$ (6)
@@
Memory buffer: Most iterative-framework-based papers generate new trajectories in each iteration and discard those from previous iterations [8, 4, 9]. Such a training manner tends to forget the infeasible region learned in early iterations. Thus, we introduce a memory replay buffer and record all the reliable infeasible data $\mathcal{R}$ identified in each iteration into the memory buffer $\mathcal{M}$. In the following iterations of learning constraints, both the current $\mathcal{R}$ and memory buffer $\mathcal{M}$ serve as infeasible data for training (see (5)).
@@ III-D Represent and Learn Policy via Constrained RL or Dynamical System Modulation
In the iterative framework discussed in the last subsection, we need to maintain a policy to generate high-reward trajectories while satisfying the currently learned constraints. We consider and compare two policies: 1) constrained RL [18], which is effective at learning constrained policies in a high-dimensional environment with complex reward function but suffers from a slow and unstable training process, and 2) dynamical system modulation (DSM) [19], which modulates the nominal policy by multiplying a rotation matrix to avoid infeasible regions. Unlike RL, the modulation requires no explicit training of the policy. Thus, DSM minimizes training time and provides a formal guarantee that the policy will never violate the learned constraints. However, DSM is applicable only for tasks where the control command is the velocity of each state , e.g., controlling the end-effector pose of a robot arm.
@@
Constrained reinforcement learning: A networked policy is trained with a modern constrained RL algorithm PID-Lagrangian[18], which offers the advantage of automatically and more stably adjusting the penalty weight. As shown in (7), it reshapes the original reward by incorporating the constraint as a penalty term into the original reward function to avoid the infeasible states. Here, $w_{p}$ is a penalty weight adjusted by PID-Lagrangian updating rules [18, 20, 21], and $c(s,a)$ is the constraint indicator function. Using this reshaped $r^{\prime}(s,a)$, a constrained optimal policy can be straightforwardly learned with the standard RL algorithm PPO [22].
$r^{\prime}(s,a)=r(s,a)-w_{p}c_{\theta}(s,a),\quad\text{where}$ (7)
$c_{\theta}(s,a)=\begin{cases}0&\text{if $\zeta_{\theta}(\bar{s})>d$ (feasible)}\\ 1&\text{if $\zeta_{\theta}(\bar{s})\leq d$ (infeasible)}\\ \end{cases}$
@@
Dynamical system modulation: DSM first learns a nominal unconstrained policy and later modulates nominal policy with the constraint network. The nominal dynamical system policy is represented as (8), where $s_{g}$ is the goal state and matrix $A(s)$ is a Gaussian Mixture Regression model learned from the known reward function and preserves stability, following [23]  [^2] In our experiment, for DSM baseline we directly set $A(s)=I$ since the task is reaching the destination in the shortest path, which is a straight line connecting the current state and the goal state..
$\pi_{n}(s)\equiv\dot{s}=A(s)(s-s_{g})$ (8)
Given a learned constraint network $\zeta_{\theta}(s)$ with state-only constraint, (9) produces a provably safe policy  [^3] The DS modulation theory guarantees that the modulated policy will never violate the learned constraint. Additionally, if the true constraint is a star-shaped obstacle and the nominal policy is linear, the modulated policy also remains globally stable [19]. $\pi(s)$ by multiplying a rotation matrix $M(s)$ constructed from $\zeta_{\theta}(s)$ following [19].
$\pi(s)=M(s)\pi_{n}(s),\quad\text{where}$ (9)
$M(s)=E(s)D(s)E^{-1}(s),$
$E(s)=[r(s),e_{1}(s),e_{2}(s),...e_{n-1}(s)],$
$n(s)=\frac{d\zeta(s)}{ds},n(s)\perp e_{1}(s)\perp e_{2}(s)\perp...e_{n-1}(s)$
$D(s)=diag[\lambda_{1}(s),\lambda_{2}(s),\lambda_{3}(s)...,\lambda_{n}(s)],$
$\lambda_{1}(s)=1-\frac{1}{\Gamma(s)},\lambda_{2,3,...,n}(s)=1+\frac{1}{\Gamma(s)},$
$\Gamma(s)=1+10*(\zeta(s)-0.5)^{0.2}$
@@
where $n(s)$ is the normal vector to the constraint boundary, $e_{i}(s)$ are the corresponding tangent vectors, $r(s)$ is the vector towards a reference point inside the obstacle, $\lambda_{i}(s)$ modifies the velocity along different directions, and $\Gamma(s)$ specifies the relative distance of current state $s$ to the constraint boundary. Intuitively, for any state $s$ close to the constraint boundary, $\Gamma(s)$ approaches 1 and $\lambda_{1}(s)$ approaches 0, resulting in zero normal velocity and preventing the agent from penetrating the obstacle.
The pseudocode of the proposed method is attached below.
@@
1:  input: Demonstration set $\mathcal{D}_{\tau}$, randomly initialized constraint network $\zeta_{\theta}$, randomly initialized policy network $\pi_{\phi}$, empty memory buffer $\mathcal{M}=\varnothing$
@@
3:     % Learning policy
4:     Evaluate current policy $\pi_{\phi}$ and adapt the penalty weight $w_{p}$ in (7) through PID-Lagrangian [18]
5:     Update the policy $\pi_{\phi}$ with the current constraint network $\zeta_{\theta}$ using constrained RL or DSM
6:     % Learning constraint with two-step PU learning
7:     Sample trajectory set $\mathcal{P}_{\tau}$ using the current policy $\pi_{\phi}$ from the same starting points as each demonstration
8:     Filter sampled set $\mathcal{P}_{\tau}$ with (6):
9:     Identify and expand reliable infeasible set $\mathcal{R}$ using (3) and (4)
10:     Update the constraint $\zeta_{\theta}$ with the demonstration $\mathcal{D}$, reliable infeasible set $\mathcal{R}$ and memory buffer $\mathcal{M}$ with (5)
11:     Save all data from $\mathcal{R}$ into the memory buffer $\mathcal{M}$
12:  until reach maximum iterations
@@ IV Experiment
IV-A Comparison of Constraint Learning Methods
To examine the performance of the proposed method, we apply it to four simulated environments to learn a 2D position constraint, a 3D position constraint, a 3D control input constraint and a constraint of an 18D cheetah robot, respectively. All codes are available at https://github.com/epfl-lasa/PUCL_learning_constraint.
@@ IV-A Comparison of Constraint Learning Methods
Baselines: Four constraint learning methods are compared here, all using constrained RL as the policy: 1) PUCL (the proposed method); 2) maximum-entropy constraint learning (MECL) [8], a popular IRL-based method that can learn a continuous arbitrary constraint function; 3) binary classifier (BC) [8], a method directly trains an NN feasibility classifier using loss (5) but with all data from $\mathcal{P}$ as infeasible data; 4) generative-PUCL (GPUCL), an alternative PUCL approach using a generative model to identify reliable infeasible data. It first learns a Gaussian Mixture Model (GMM) of the expert distribution, and then identifies reliable infeasible data as the data from $\mathcal{P}$ with the lowest likelihood of being generated by the GMM. GPUCL is also first formulated in this paper but we do not discuss in detail due to its weaker performance.
Metrics: we present the evaluation utilizing two metrics: 1) the IoU (the intersection over union) index, which measures the correctness of the learned constraints. We uniformly sample points in the state space, and compute IoU as (Number of points which are both predicted to be infeasible and truly infeasible) divided by (Number of points either predicted to be infeasible or truly infeasible); the constraint models of all four methods use 0.5 as the decision threshold. 2) the unsafe rate, the per-step true constraint violation rate of the policy learned from the learned constraint.
2D/3D Position Constraints: We first study the two obstacle-avoidance constraints of the end-effector of a robot arm, one in 2D position space and another in 3D position space. We defined a goal-reaching task, where the agent is initialized randomly and rewarded to reach a goal state while staying away from the unknown obstacles. The state space consists of 2D/3D coordinates of the end-effector and the control variables are its corresponding 2D/3D velocities, capped at 0.58m/s. The reward functions in both tasks are the negative of the Cartesian path distance to reach the goal, a natural choice for a goal-reaching task. The constraint networks in both tasks take only the state as input, i.e., $\bar{s}=s$.
In the 2D reaching task, the true constraint involves an irregular obstacle composed of two ellipses as shown in Fig. 4. In the 3D reaching task, the true constraint is avoiding the tall cylinders. We stress that learning such constraints in our setting is nontrivial and has not been well addressed by prior works. Specifically, our setting does not assume prior knowledge of the obstacles size, shape, or location, while relevant studies [13] typically require a known parameterization of obstacles and only infer a handful of parameters (e.g., height and radius of cylinders). Furthermore, the studied constraints are nonlinear, not just boxes or hyperplanes, which contrasts with previous works that only learn a plain constraint of a single dimension such as coordinate $x\geq-3$ [8, 9]. Finally, unlike [12], our method does not require a differentiable environmental model.
The expert demonstration sets for 2D and 3D tasks consist of 4 and 10 safe trajectories, respectively. All demonstrations are generated by an entropy-regularized suboptimal RL agent, trained assuming full knowledge of the true constraint.
@@
A list of important hyper-parameters of our algorithm is given in Table I. For the neural network architectures, the number of hidden units in each layer is mentioned. The hidden activation functions in all networks are Leaky ReLU. The output activation function of the constraint network is the sigmoid function, whereas that of the policy and value network is tanh. Other baselines all adopt identical network architecture.
TABLE I: List of hyper-parameters of PUCL.
@@
Distance metric Euclidean Euclidean
$\mathcal{R}$ distance threshold $d_{r}$ 0.12 0.03
@@
The IoU index and the unsafe rate in the two tasks are presented in Fig. 5. Each group of 10 runs uses the same demonstrations and hyper-parameters, but differ in the initialization parameters for the networks. The proposed method PUCL exhibits superior performance compared to the baseline across both environments and metrics. It achieves a near-zero unsafe rate at the convergence of the training in both environments and outperforms baseline MECL. GPUCL adopts a similar framework to PUCL but performs worse than PUCL, possibly due to the difficulty of fitting a good expert distribution with a very limited numbers of demonstrations.
@@
To provide readers with an intuitive understanding, we include visualizations of the 2D constraint learned with PUCL and MECL in Fig. 4. Comparing the learned constraints (red region) with the true constraints (white ellipses) confirms the effectiveness of our method in acquiring a model of a nonlinear constraint. The baseline method MECL learns a less accurate constraint, and even some states visited by the demonstrations are classified as infeasible.
While our method is effective, it can have a relatively strong dependency on the distance threshold $d_{r}$. To clearly illustrate the impact of this hyperparameter on performance, we present a sensitivity analysis by varying $d_{r}$ in the 3D task (see Fig. 3).
Figure 3: Sensitivity analysis of the distance threshold $d_{r}$ on performance. The left y-axis represents the unsafe rate (lower is better), and the right y-axis represents the IoU (higher is better). The curve is averaged on 5 runs with different random seeds.
(a) PUCL (ours)
(b) MECL
Figure 4: True constraint and constraint learned with the proposed method PUCL and baseline MECL [8] in the 2D reaching task. The two white ellipses represent the boundary of the true constraint. The heat map is a visualization of the learned constraint $\zeta_{\theta}$, where the red region of $\zeta=0$ represents infeasibility, while the blue regions of $\zeta=1$ indicate feasibility. The trajectories of the demonstrations and the policy are shown in green and yellow, respectively. The red points correspond to identified reliable infeasible data $\mathcal{R}\cup\mathcal{M}$ (only for PUCL method).
Figure 5: The IoU index (higher is better) and unsafe rate (lower is better) in three environments (top: 2D reaching, middle: 3D reaching, bottom: blocked half cheetah). The x-axis indicates the training process, the number of timesteps the agent takes in the environment. All the results are the average of 10 independent runs, and the shaded area represents one standard deviation.
Control Input Constraint: To demonstrate our method’s ability to learn control input constraint, we modify the 3D reaching task by imposing stricter maximum velocity constraints. The original permitted velocities limits are $|v_{x}|,|v_{y}|,|v_{z}|\leq 0.58m/s$. Imagine that when the human users demonstrate the motion, they may expect the robot to move slower according to their personal preference, e.g., a stricter velocity constraint $|v_{x}|\leq 0.48m/s,|v_{y}|\leq 0.58m/s,|v_{z}|\leq 0.19m/s$. The agent is tasked with inferring this user-specific velocity constraint from demonstrations. Importantly, although the true constraints follow a box form, the agent is not given any prior knowledge of this structure; it cannot assume a box constraint and simply find maximum values from the data. For simplicity, the obstacle-avoidance constraint is assumed to be known in this task, and only the velocity constraint is learned, thus $\bar{s}=a$.
We collected 30 demonstrations from an expert RL agent and applied PUCL and MECL to learn the velocity constraint. For PUCL, we used the Euclidean distance with a threshold of $d_{r}=0.01$. The results, averaged over 8 runs, showed that PUCL achieved an IoU score of $0.87\pm 0.08$, compared to $0.77\pm 0.26$ for MECL. In terms of unsafe rate, PUCL achieves an unsafe rate of only $(3.47\pm 3.26)\%$ compared to $(28.1\pm 19.2)\%$ for MECL. Furthermore, the velocity distribution output by the policy learned with PUCL is shown in Fig. 6, demonstrating that almost all data points fall within the true safe range.
Figure 6: Velocity distribution of the policy learned with PUCL (Left: $v_{x}$, Right: $v_{z}$). The horizontal axis represents the velocity, and the red line represents the true constraint threshold.
Constraint in high-dimensional space: The previous tasks extracts constraints only from 2D/3D spaces. To showcase our method extends to high-dimensional, heterogeneous state space, we apply it to an 18D blocked half cheetah environment introduced by baseline papers [8, 9]. The agent controls a cheetah robot with two legs by applying torques on each joint motors. The reward is determined by the distance it walks between the current and the previous time step. The constraint blocks the region with x-coordinate $\leq-3$, so the robot is only allowed to move in the region with x-coordinate between $-3$ and $\infty$. The input to the constraint network is 18-dimensional state space, including positional and angular values of different body parts of the cheetah, followed by their velocities, making constraint extraction challenging due to this high-dimensional space.
We collected 30 demonstrations and run PUCL to learn constraint. Since each dimension of the state space represents distinct physical properties and scales, we identify the reliable infeasible data based on the standardized dataset as discussed in III-B. Euclidean distance is used as the distance measure  [^4] We also tested Manhattan distance and Chebyshev distance. Both achieves similar performance with appropriately selected distance thresholds. with a threshold  [^5] This value is tuned following the discussion in III-B for heterogeneous cases: we first collect $\mathcal{P}$ using an unconstrained policy and compute the score $\mathcal{S}_{\mathcal{D}}(\bar{s})$ of each state from $\mathcal{P}$. We hypothesis that around 80% of states in $\mathcal{P}$ are reliable infeasible data, and this corresponds to a threshold of 2.6. $d_{r}=2.6$. The learning performance is presented in Fig. 5, where PUCL achieves the highest IoU metrics and the second best unsafe rate. This confirms the ability of PUCL for learning constraints in high-dimensional state spaces.
IV-B Comparison of Policy Representation and Learning Methods